OpenStudy (anonymous):

Find the slope of each line. 12x - 15y =37 would I divide right? or do I still do the rise over run technique?

OpenStudy (anonymous):

4/5

OpenStudy (anonymous):

isolate y so 12x - 37 = 15y 12/15 x - 37/15 = y 4/5 is the slope

hero (hero):

Moneybird, define "infinite slope".

hero (hero):

A vertical line has infinite slope and infinite slope is defined, therefore, a vertical line x = 5 is defined. And therefore, the notion that the slope of a vertical line is undefined is false.

OpenStudy (anonymous):

But the denominator is zero when you plug it into the slope formula

hero (hero):

\[\frac{1}{0} = \infty , \frac{1}{\infty} = 0\] \[0*\infty = 1, \infty * 0 = 1\] \[0*\infty = \infty * 0\] \[1 = 1\]

OpenStudy (anonymous):

I agree but still infinite slope is undefined

hero (hero):

Infinity means undefined only in the sense that there is no finite number to describe it

OpenStudy (anonymous):

There is no finite number to describe infinity since it only means a very big number

hero (hero):

Infinity exists at every point so every number is infinite and all numbers are undefined. The real number line is an illusion created by man's attempt to quantify it

OpenStudy (anonymous):

No infinity doesn't mean every number. infinity means that a number is too big and without end or bound.

hero (hero):

Every number is a component of infinity. Every number is part of the set of infinite numbers. The real number line is a visual representation of infinity.

hero (hero):

Define the number six

OpenStudy (anonymous):

6 is a real and finite number

hero (hero):

You can't define the number six without defining 1's and 0's, both components of infinity with zero itself being the origin, a point at infinity. A number is a basic constituent of infinity.

hero (hero):

That's not a definition of six because you can say the same thing about 7 or any other number. You have to define six and six only

OpenStudy (anonymous):

But it's still undefined for the slope of a vertical line. Because it could be -infinity or infinity slope

hero (hero):

The fact that x = -6 means "infinite slope" proves that all numbers are infinite. All numbers are connected with infinity

hero (hero):

On that note, I'm going to bed. Goodnight.

OpenStudy (anonymous):

Same here good night